Can you please help me and identifies if it is one solution,no solution or infinte solution:
3x+y=1
2x-y=-6

Answer by rapaljer(4671)

Just add these two equations together:
3x+y=1
2x-y=-6

5x=-5
x=-1

Substitute back into the first equation:
3x+y=1
3(-1)+Y=1
-3+y=1
y=4

Check in the second equation:
2x-y=-6
2(-1)-(4)=-6
-2-4=-6 It checks!!

(-1, 4)


For an easier explanation of this topic, please see my website by clicking on my tutor name "rapaljer" anywhere in algebra.com. On my homepage, look for "Basic, Intermediate, and College Algebra: One Step at a Time." Select "Basic Algebra" and look for the appropriate sections in Chapter 4 of my Basic Algebra book. It's all FREE, just like algebra.com!!

R^2